# On the kissing number of the cross-polytope

### v2

@@ 3 The lattice kissing number of $K_{n}$
In this section, we focus on the lattice kissing number problem for $K_{n}$, and prove the following theorem:
@@
Let $L\subset\mathbb{R}^{4}$ be a lattice achieving $\kappa_{L}(K_{4})$, and let $\mathcal{M}$ be a set containing exactly one of each $\pm$ pairs of vectors in $\mathcal{M}_{K_{4}}(L)$. Suppose that $|\mathcal{M}|\geq 20$. Note that if $x,y\in\mathcal{M}$ are distinct vectors such that $x\equiv y\pmod{2L}$, then $\frac{x\pm y}{2}\in\mathcal{M}_{K_{4}}(L)$ and the triangle inequality implies that $\text{supp}(x)\cap\text{supp}(y)=\emptyset$, and in particular $|\text{supp}(x)|\leq 2$ or $|\text{supp}(y)|\leq 2$.
Since there are $152^{4}-1=$ 15 non-zero classes in $L/2L$, the following is true: $\mathcal{M}$ contains $k\in\{0,1,2,3,4\}$ vectors whose support has size one and $5-k$ pairs $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$ of vectors such that $x_{1}^{i}\equiv x_{2}^{i}\pmod{2L}$ and $|\text{supp}(x_{1}^{i})|=|\text{supp}(x_{2}^{i})|=2$ for all $i=1,\dots,5-k$. Note that there must exist at least one pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ satisfying $\lVert x_{1}^{i_{1}}\rVert_{\infty}=\lVert x_{2}^{i_{1}}\rVert_{\infty}=\frac{1}{2}$. Indeed, if $x,y\in\mathcal{M}$ share a max-coordinate, then $\lVert x\rVert_{\infty}\leq\frac{1}{2}$ or $\lVert y\rVert_{\infty}\leq\frac{1}{2}$, and every $x\in\mathcal{M}$ with $|\text{supp}(x)|>1$ has $|x_{i}|\leq\frac{1}{2}$ for all $i\in\cup_{x\in\mathcal{M}:|\text{supp}(x)|=1}\text{supp}(x)$. If $k\leq 2$, then in addition to the pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ there are two vectors $x_{3},x_{4}\in\mathcal{M}$ such that $\lVert x_{3}\rVert_{\infty}=\lVert x_{4}\rVert_{\infty}=\frac{1}{2}$ and $|\text{supp}(x_{3})|=|\text{supp}(x_{4})|=2$. Thus, replacing $L$ by a $K_{4}$-equivalent lattice, we may assume that $L$ contains the lattice $L^{\prime}$ generated by $\mathbb{Z}^{4}$, the vectors $(\frac{1}{2},\frac{1}{2},0,0)$, $(0,0,\frac{1}{2},\frac{1}{2})$ and their midpoint $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$.
Since $|\mathcal{M}_{K_{4}}(L^{\prime})|=20$ and $\rho(L^{\prime})\leq 1$ by the next lemma, $L$ contains in addition to $L^{\prime}$ at least one of its deep holes. By the next lemma, $L\setminus L^{\prime}$ contains a vector of one of the following shapes.
@@
for all $0\leq x_{1},x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{1},x_{2}\in\{0,\frac{1}{2}\}$, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ and so $L$ contains $\frac{1}{2}D_{4}^{+}$.
@@
for all $\frac{1}{4}\leq x_{1}\leq\frac{1}{2}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Similarly,
@@
for all $0\leq x_{1}\leq\frac{1}{4}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Thus, $L$ contains $(x_{1},x_{1}-\frac{1}{2},0,\frac{1}{2})$ or $(x_{1},x_{1}-\frac{1}{2},\frac{1}{2},0)$ so it contains $(2x_{1}-\frac{1}{2},2x_{1}-\frac{1}{2},0,0)$. Since $-\frac{1}{2}\leq 2x_{1}-\frac{1}{2}\leq\frac{1}{2}$, we must have $x_{1}\in\{0,\frac{1}{4},\frac{1}{2}\}$. Thus, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ or $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$ so it contains $\frac{1}{2}D_{4}^{+}$ or the lattice $L_{0}$ spanned by $L^{\prime}$ and $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$. The assumption $v_{2}^{\prime}\in L$ leads to the same conclusion.
@@
In conclusion, $L$ contains $\frac{1}{2}D_{4}^{+}$, $L_{0}$, $L_{1}$ or $L_{1}^{\prime}$. Suppose that $L$ strictly contains one of these lattices. Then by the reasoning above, $L$ contains at least two of these lattices. This however implies that $\min_{x\in L\setminus\{0\}}\lVert x\rVert_{1}<1$, a contradiction. In conclusion, $L\in\{\frac{1}{2}D_{4}^{+},L_{0},L_{1},L_{1}^{\prime}\}$. We check that
$|\mathcal{M}_{K_{4}}(\frac{1}{2}D_{4}^{+})|=40,\quad|\mathcal{M}_{K_{4}}(L_{0})|=36,\quad|\mathcal{M}_{K_{4}}(L_{1})|=|\mathcal{M}_{K_{4}}(L_{1}^{\prime})|=28,$
and conclude that $L=\frac{1}{2}D_{4}^{+}$.
@@
The deep holes of a direct sum $L_{1}\oplus L_{2}$ of lattices $L_{1}$, $L_{2}$ are given by $(x,y)$, where $x$ is a deep hole of $L_{1}$ and $y$ is a deep hole of $L_{2}$. Thus, it suffices to show that the deep holes of $\mathcal{H}_{2}$ consist of the translates by vectors in $\mathcal{H}_{2}$ of the line segments $\{(x,\pm(\frac{1}{2}-x)):0\leq x\leq\frac{1}{2}\}$. This is clear. ∎
@@ 4.1 An asymptotic upper bound for $\kappa_{T}(K_{n})$
In our estimations, the following limit will appear often: for a fixed real number $0\leq a\leq 1$, $\lim_{n\to\infty}\frac{1}{n}\log_{2}\binom{n}{\lfloor an\rfloor}=H\left(a\right)$, where $H(s)=-s\log_{2}(s)-(1-s)\log_{2}(1-s)$ and $H(0)=0=H(1)$ is the binary entropy function defined for $0\leq s\leq 1$.
@@
Let $X=\{x_{1},\dots,x_{N}\}$ be a cross-polytope kissing configuration, so that the interiors of $x_{i}+\frac{1}{2}K_{n}$ and $x_{j}+\frac{1}{2}K_{n}$ are disjoint for all $i\neq j$. Let $X^{\prime}=\{x\in X:\text{$x$ has at least $\lfloor cn\rfloor$ coordinates of absolute value}\geq\frac{b}{n}\}$. We select parameter values $c=0.296$ and $b=0.334$. Note that for $x\in X\setminus X^{\prime}$, the sum of the absolute values of $n-\lfloor cn\rfloor$ smallest coordinates (in absolute value) is at most $b(1-c)+\frac{b}{n}=a$. Thus, if $x,x^{\prime}\in X\setminus X^{\prime}$ are such that the positions of $\lfloor cn\rfloor$ largest coordinates coincide (say positions $i=1,\dots,\lfloor cn\rfloor$), then $\sum_{i=1}^{\lfloor cn\rfloor}|x_{i}-x_{i}^{\prime}|\geq 1-2a$. Hence, the interiors of the $\lfloor cn\rfloor$-dimensional bodies $(x_{1},\dots,x_{\lfloor cn\rfloor})+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ and $(x_{1}^{\prime},\dots,x_{\lfloor cn\rfloor}^{\prime})+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ are disjoint, and both are contained in $\frac{3-2a}{2}K_{\lfloor cn\rfloor}$. By a volume comparison, we obtain
@@
from which it follows that $M\leq 2.91616^{(1+o(1))n}$. ∎
@@ 4.2 An asymptotic lower bound for $\kappa_{T}(K_{n})$
Let $m>0$ and let $\mathcal{S}$ be a finite subset of $\partial(mK_{n})$ such that for all $x\in\mathcal{S}$, the cardinality of the set $\mathcal{B}(x)=\{y\in\mathcal{S}:\lVert x-y\rVert_{1}<m\}$ is independent of the center $x$. This is achieved when $\mathcal{S}$ is the set of signed permutations of one vector $v\in\mathbb{R}^{n}$. Then by the union bound,
@@
of cardinality $|\mathcal{X}_{m_{1},m_{2}}|=\binom{n}{m_{1}}\binom{n-m_{1}}{m_{2}}2^{m_{1}+m_{2}}$. Note that $\lVert x\rVert_{1}=m_{1}+2m_{2}:=m$ for all $x\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let $\mathcal{B}_{m_{1},m_{2}}(x)=\{y\in\mathcal{X}_{m_{1},m_{2}}:\lVert x-y\rVert_{1}<m\}$. Clearly $|\mathcal{B}_{m_{1},m_{2}}(x)|$ does not depend on the center $x$ so we may consider the ball centered at $x_{0}=(1,\dots,1,2,\dots,2,0,\dots,0)\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let

### v3

@@ Abstract
Key words. Kissing number, Cross-polytope, Lattice.
MSC Codes. 52C17, 11H31.
@@ 3 The lattice kissing number of $K_{n}$
In this section, we focus on the lattice kissing number problem for $K_{n}$, and prove the following theorem.
@@
Let $L\subset\mathbb{R}^{4}$ be a lattice and let $\mathcal{M}$ be a set containing exactly one of each $\pm$ pairs of vectors in $\mathcal{M}_{K_{4}}(L)$. Suppose that $|\mathcal{M}|\geq 20$. Note that if $x,y\in\mathcal{M}$ are distinct vectors such that $x\equiv y\pmod{2L}$, then $\frac{x\pm y}{2}\in\mathcal{M}_{K_{4}}(L)$ and the triangle inequality implies that $\text{supp}(x)\cap\text{supp}(y)=\emptyset$, and in particular $|\text{supp}(x)|\leq 2$ or $|\text{supp}(y)|\leq 2$.
Since there are $152^{4}-1=$ 15 non-zero classes in $L/2L$, the following is true: $\mathcal{M}$ contains $k\in\{0,1,2,3,4\}$ vectors whose support has size one and $5-k$ pairs $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$ of vectors such that $x_{1}^{i}\equiv x_{2}^{i}\pmod{2L}$ and $|\text{supp}(x_{1}^{i})|=|\text{supp}(x_{2}^{i})|=2$ for all $i=1,\dots,5-k$. If $x,y\in\mathcal{M}$ are distinct vectors sharing a max-coordinate, then $\lVert x\rVert_{\infty}\leq\frac{1}{2}$ or $\lVert y\rVert_{\infty}\leq\frac{1}{2}$. In particular, if $|\text{supp}(x)|\geq 2$, then $|x_{i}|\leq\frac{1}{2}$ for all $i\in\cup_{x\in\mathcal{M}:|\text{supp}(x)|=1}\text{supp}(x)$. Then by the pigeonhole principle, among the $2(5-k)$ vectors $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$, there are at most $4-k$ with max-norm $>\frac{1}{2}$. In particular, there exists a pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ satisfying $\lVert x_{1}^{i_{1}}\rVert_{\infty}=\lVert x_{2}^{i_{1}}\rVert_{\infty}=\frac{1}{2}$, and moreover, $\mathcal{M}$ contains in addition to the pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ at least $4-k$ vectors $\{y_{i}\}_{i=1}^{4-k}$ with $\lVert y_{i}\rVert_{\infty}=\frac{1}{2}$ and $|\text{supp}(y_{i})|=2$ for all $i=1,\dots,4-k$. Recall that $\text{supp}(x_{1}^{i_{1}})\cap\text{supp}(x_{2}^{i_{1}})=\emptyset$. Thus, replacing $L$ by a $K_{4}$-equivalent lattice, we may suppose that $x_{1}^{i_{1}}=(\frac{1}{2},\frac{1}{2},0,0)$ and $x_{2}^{i_{1}}=(0,0,\frac{1}{2},\frac{1}{2})$. The vectors $x_{1}^{i_{1}}$, $x_{2}^{i_{1}}$, $\frac{x_{1}^{i_{1}}+x_{2}^{i_{1}}}{2}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ together with the vectors $\{y_{i}\}_{i=1}^{4-k}$ and $\{x\in\mathcal{M}:|\text{supp}(x)|=1\}$ span a lattice which contains the lattice $L^{\prime}$ generated by $\mathbb{Z}^{4}$, $(\frac{1}{2},\frac{1}{2},0,0)$, $(0,0,\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$.
Since $|\mathcal{M}_{K_{4}}(L^{\prime})|=20$, the containment $L^{\prime}\subsetneq L$ is strict. Since $\rho(L^{\prime})\leq 1$ by the next lemma, $L$ contains a deep hole of $L^{\prime}$, which is also a deep hole of the lattice $\mathcal{H}_{2}\oplus\mathcal{H}_{2}$ generated by $\mathbb{Z}^{4}$, $(\frac{1}{2},\frac{1}{2},0,0)$ and $(0,0,\frac{1}{2},\frac{1}{2})$. Thus, by the next lemma, $L\setminus L^{\prime}$ contains a vector of one of the following types.
@@
for all $0\leq x_{1},x_{2}\leq\frac{1}{2}$, with equality if and only if $x_{1},x_{2}\in\{0,\frac{1}{2}\}$. Hence, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$. The lattice $\frac{1}{2}D_{4}^{+}$ is generated by $L^{\prime}$ and $(\frac{1}{2},0,\frac{1}{2},0)$. We conclude that $\frac{1}{2}D_{4}^{+}\subseteq L$.
@@
for all $\frac{1}{4}\leq x_{1}\leq\frac{1}{2}$, $0\leq x_{2}\leq\frac{1}{2}$, with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Similarly,
@@
for all $0\leq x_{1}\leq\frac{1}{4}$, $0\leq x_{2}\leq\frac{1}{2}$, with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Thus, $v_{2}=(x_{1},x_{1}-\frac{1}{2},0,\frac{1}{2})$ or $v_{2}=(x_{1},x_{1}-\frac{1}{2},\frac{1}{2},0)$. This implies that $L$ contains $(2x_{1}-\frac{1}{2},2x_{1}-\frac{1}{2},0,0)$. Since $-\frac{1}{2}\leq 2x_{1}-\frac{1}{2}\leq\frac{1}{2}$, we must have $x_{1}\in\{0,\frac{1}{4},\frac{1}{2}\}$. Thus, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ or $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$, so it contains $\frac{1}{2}D_{4}^{+}$ or the lattice $L_{0}$ spanned by $L^{\prime}$ and $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$. The assumption $v_{2}^{\prime}\in L$ leads to the same conclusion.
@@
In conclusion, $L$ contains $\frac{1}{2}D_{4}^{+}$, $L_{0}$, $L_{1}$ or $L_{1}^{\prime}$. If the containment is strict, then $L$ contains an additional deep hole as above, which implies that $L$ contains two of these lattices. This however implies that $\min_{x\in L\setminus\{0\}}\lVert x\rVert_{1}<1$, a contradiction. In conclusion, $L\in\{\frac{1}{2}D_{4}^{+},L_{0},L_{1},L_{1}^{\prime}\}$. We check that
$|\mathcal{M}_{K_{4}}(\frac{1}{2}D_{4}^{+})|=40,\quad|\mathcal{M}_{K_{4}}(L_{0})|=36,\quad|\mathcal{M}_{K_{4}}(L_{1})|=|\mathcal{M}_{K_{4}}(L_{1}^{\prime})|=28.$
We are left with only one possibility: $L=\frac{1}{2}D_{4}^{+}$.
@@
The deep holes of a direct sum $L_{1}\oplus L_{2}$ of lattices $L_{1}$, $L_{2}$ are given by $(x,y)$, where $x$ is a deep hole of $L_{1}$ and $y$ is a deep hole of $L_{2}$. Thus, it suffices to determine the deep holes of $\mathcal{H}_{2}$. Note that $\mathcal{H}_{2}$ is a scaled rotation of the lattice $\mathbb{Z}^{2}$, and $K_{2}$ is a scaled rotation, with the same factor, of the cube $Q_{2}=\{x\in\mathbb{R}^{2}:\lVert x\rVert_{\infty}\leq 1\}$. The deep holes of $\mathbb{Z}^{2}$ with respect to the norm $\lVert\cdot\rVert_{\infty}$ are the sets of vertical and horizontal lines passing through $(\frac{1}{2},\frac{1}{2})+\mathbb{Z}^{2}$. Now rotating back to the original situation, we see that the deep holes of $\mathcal{H}_{2}$ consist of the translates by vectors in $\mathcal{H}_{2}$ of the line segments $\{(x,\pm(\frac{1}{2}-x)):0\leq x\leq\frac{1}{2}\}$. ∎
@@ 4.1 An asymptotic upper bound for $\kappa_{T}(K_{n})$
In our estimations, the following limit will appear often: for a fixed real number $0\leq a\leq 1$, $\lim_{n\to\infty}\frac{1}{n}\log_{2}\binom{n}{\lfloor an\rfloor}=H(a)$, where $H(s)=-s\log_{2}(s)-(1-s)\log_{2}(1-s)$ and $H(0)=0=H(1)$ is the binary entropy function defined for $0\leq s\leq 1$.
@@
Let $X=\{x_{1},\dots,x_{N}\}$ be a cross-polytope kissing configuration, so that the interiors of $x_{i}+\frac{1}{2}K_{n}$ and $x_{j}+\frac{1}{2}K_{n}$ are disjoint for all $i\neq j$. For $x\in X$, define $I_{x}=\{i:|x_{i}|\geq\frac{b}{n}\}\subseteq\{1,\dots,n\}$ and let $X^{\prime}=\{x\in X:|I_{x}|\geq\lfloor cn\rfloor\}$. We select parameter values $b=0.334$ and $c=0.296$. For each $x\in X\setminus X^{\prime}$, assign a set $S_{x}\subseteq\{1,\dots,n\}\setminus I_{x}$ such that $|S_{x}|=n-\lfloor cn\rfloor$. Then $\sum_{i\in S_{x}}|x_{i}|\leq b(1-c)+\frac{b}{n}=a$. Thus, if $x,x^{\prime}\in X\setminus X^{\prime}$ are distinct points such that the sets $S_{x}$, $S_{x^{\prime}}$ coincide, then $\sum_{i\not\in S_{x}}|x_{i}-x_{i}^{\prime}|\geq 1-2a$, since otherwise $\sum_{i=1}^{n}|x_{i}-x_{i}^{\prime}|\leq\sum_{i\not\in S_{x}}|x_{i}-x_{i}^{\prime}|+\sum_{i\in S_{x}}(|x_{i}|+|x_{i}^{\prime}|)<1-2a+2a=1$, a contradiction. Hence, the interiors of the $\lfloor cn\rfloor$-dimensional bodies $(x_{i})_{i\not\in S_{x}}+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ and $(x_{i}^{\prime})_{i\not\in S_{x^{\prime}}}+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ are disjoint, and both are contained in $\frac{3-2a}{2}K_{\lfloor cn\rfloor}$. By a volume comparison, we obtain
@@
from which it follows that $M\leq 2.91616^{(1+o(1))n}$, by routine estimation. ∎
@@ 4.2 An asymptotic lower bound for $\kappa_{T}(K_{n})$
Let $m>0$ and let $\mathcal{S}$ be a finite subset of the boundary $\partial(mK_{n})$ such that for all $x\in\mathcal{S}$, the cardinality of the set $\mathcal{B}(x)=\{y\in\mathcal{S}:\lVert x-y\rVert_{1}<m\}$ is independent of the center $x$. This is achieved when $\mathcal{S}$ is the set of signed permutations of one vector $v\in\mathbb{R}^{n}$. Then by the union bound,
@@
of cardinality $|\mathcal{X}_{m_{1},m_{2}}|=\binom{n}{m_{1}}\binom{n-m_{1}}{m_{2}}2^{m_{1}+m_{2}}$. Note that $\lVert x\rVert_{1}=m_{1}+2m_{2}:=m$ for all $x\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let $\mathcal{B}_{m_{1},m_{2}}(x)=\{y\in\mathcal{X}_{m_{1},m_{2}}:\lVert x-y\rVert_{1}<m\}$. Clearly, $|\mathcal{B}_{m_{1},m_{2}}(x)|$ does not depend on the center $x$ so we may consider the ball centered at $x_{0}=(1,\dots,1,2,\dots,2,0,\dots,0)\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let